Search
Date | Author | Entry | File | Actions |
---|---|---|---|---|
2024-11-20 | kahmark |
Lux marginal land site visited by Kevin/Stacey today. Found many corner posts and sockets pulled out and also north side of G10R3 found soil moisture probes dug up and chewed. Will not be able to repair. They are soil moisture profiler sensors for this plot.. |
|
|
2024-11-20 | dave |
BCSE202410071104S16S17BGC run had some submitted samples out of order. All samples run on GC in numerical order. Series 16 samples out of order, submitted in following sequence: …10, 11, 16, 13, 14, 15, 20, 17, 18, 19, 12, check, 21, 22…. |
|
|
2024-11-08 | kahmark |
KK/SV finished switchgrass gradient deep cores today, 6 plots (12 cores). Total of 64, two per plot. Core 1 was always in the south half of the plot and Core 2 in the north. |
|
|
2024-11-06 | kahmark |
Completed all but 12 deep cores on the Switchgrass Gradient project. KK/DW Still need to collect inorganic N samples and another gas sample due soon. |
|
|
2024-10-30 | kahmark |
Removed all chambers from SWFG study for harvest. |
|
|
2024-10-30 | kahmark |
Deep core of Switchgrass N Rate trial started today. KK/DW We cored the western 6 plots, 12 cores just after strip harvest. |
|
|
2024-10-25 | kahmark |
Adjusted all chamber heights in gashog to reflect reinstalled ht. |
|
|
2024-10-24 | dave |
GC Run SWFG series39 10/8/2024 2 vials out of order at sampling: 14 and 15 were switched at sampling, so will be run in the order …12, 13, 15, 14, 16, 17…. GC Run SWFG series41 10/23/2024 2 vials out of order at sampling: 93 and 94 were switched at sampling, so will be run in the order …92, 94, 93, 95, 96…. |
|
|
2024-10-24 | kahmark |
Removed G5 autochambers for microplot harvest in the afternoon. Turned off chambers on gashog. |
|
|
2024-10-23 | sven |
Dave and Kevin replaced the chambers on G1 and restarted sampling. G1-3 overextending. Raised all tubing off the ground. Chambers removed 10/22 pm |
|
|
2024-09-24 | dave |
Checked all chambers in G1 and G5, except G5-5 (closed at the time). Trimmed corn leaves out of the path of the closures on G1 chambers. Most G5s were in good shape and required minimal trimming on this visit. All G5 lid connections appeared to be intact and working properly. |
|
|
2024-08-16 | dave |
BCSE202407190731S12S13BGC: All vials run in consecutive order. Vials submitted by techs in this order for Series 12: 1, 18, 3, 4, 5, 14, 7, 8…. ….12, 13, 6, 15, 16, 17, 2, 19, 20, 21…. all other vials in order. |
|
|
2024-07-15 | dave |
G5-3 chamber lid fixed. May have been okay up to point of checking lids, but arm attachment popped off when checking. Fixed. |
|
|
2024-07-03 | sven |
I removed the radio, datalogger, TDR100, relay, battery and antenna from the DF site. |
|
|
2024-06-25 | dave |
Running 3 series of GLBRC gas samples. Notes on Series 6 from Techs - Vials 1-40 swapped, i.e. vials 1-20 are actually vials 21-40 regarding site sampling. Series 7 no notes. Series 8 - a number of vials out of order. All samples run in numerical order. Series 8 vial order in trays: 1-40 ok, then 41, 42, 59, 60, 45, 46, 55, 56, 49, 50, 51, 52, 53, 54, 47, 48, 57, 58, 43, 44, 61, 62, 63,… |
|
|
2024-06-20 | kahmark |
Removing G1 lids for herbicide today. 4 Hr REI. |
|
|
2024-06-17 | dave |
Approximately 4pm, trimmed switchgrass and any other plant material growing out of the G5 automated chambers. Also checked lids and trimmed any plants outside of chambers that were interfering with lid seals. |
|
|
2024-06-05 | dave |
GLBRC Series 4 vials. Note from techs that vials 1 and 21 were swapped (i.e. 21, 2, 3, 4, 5…20, 1, 22, 23, etc.). Still run on GC in consecutive order (1, 2, 3, etc…) |
|
|
2024-05-30 | kahmark |
It appears the ICOS needs to have the mirrors cleaned. Likely some water entered the system over the weekend. Caliubration worked but sensitivity is down, and noise up. |
|
|
2024-05-29 | kahmark |
Trailer went down over the weekend. ICOS concentrations way off and flow blocked. Found issue with valves, then recalibrated ICOS with our reference gas. Added lids back on G1 last Friday. Currently not running trailer until morning. |
|
|
2024-05-21 | kahmark |
Removed chambers on the SWFG for fertilization. Will install tomorrow and clip in AM, then sample after lunch. |
|
|
2024-05-20 | kahmark |
G1-6 showing negative CO2. Will check. Also removing chambers on SWFG for fertilization today. |
|
|
2024-05-20 | dave |
Moved chamber 5-3 approximately 1 meter to the west to encompass some switchgrass plants, approximately 2pm. Trimmed all chambers in G5. Replaced clamp head on G1-5 (previous one had popped off a couple times). |
|
|
2024-05-20 | dave |
G5-3 updated chamber depths after moving: corner heights - 32.5cm, 32.2cm, 31.0cm, 30.0cm; center height - 29.5cm. |
|
|
2024-05-15 | dave |
BCSE automated chamber notes: G1-5: discovered a missing nut/bolt combo on lid that may leak air. Plugged temporarily with weatherstripping on both sides of hole. G5-1, 5-3, 5-4 chambers have very little to no switchgrass growth inside the chambers. May be time to move them to be more representative of the plot. G5-2: switchgrass growth to a few inches above top of chamber. This was trimmed back to being even/slightly below top of chamber. |
|
|
2024-05-13 | dave |
G1 chambers flagged and removed for planting and liming today. Chambers all removed by 9:30am. Will replace when farm management notifies us they have completed their work. |
|
|
2024-05-13 | dave |
Re-installed G1 chambers after planting and liming management earlier in the day. Chambers all had to be shifted somewhat east to line up between newly planted rows. Four corner average heights recorded, along with one center measurement. (All numbers in centimeters.) G1-1: 30.7, 30.5, 29.7, 30.1cm, 30.5 center G1-3: 15.5, 15.6, 15.5, 15.6cm, 15.7cm center G1-2: 29.8, 30.1, 28.4, 28cm, 29cm center G1-6: 29.2, 28.9, 29.3, 29.5cm, 30cm center G1-5: 30, 29.8, 29.2, 28.9cm, 29.5cm center G1-4: 29.3, 29.5, 27.8, 28.2cm, 29.5cm center |
|
|
2024-02-05 | dave |
Checked automated chambers. No issues seen with 1-6, though reports of issues from remote data. 5-4 was missing a clamp head. Found it and replaced it. May need a new one as a long-term fix. |
|
|
2024-01-17 | kahmark |
Disconnected the incoming and outgoing sampling lines from the Licor and ICOS in the trailer until we can clear the chambers and get them operational. |
|
|
2024-01-11 | kahmark |
G5-6, G1-3, G1-5 all overextended. Turned down compressor outgoing pressure to just over 52psi. Will shut down system today prior to snowstorm and incoming polar vortex. |
|
|
2023-12-21 | dave |
Found chamber G5-6 (control) had a hyperextended lid on the east side. Put back in place. Going back out to see if I can adjust it with a screwdriver to prevent it from recurring. |
|
|
2023-10-29 | vanderws |
Process BCSE deep cores - Sidney, Alex |
|
|
2023-10-28 | vanderws |
Process BCSE deep cores - Sidney, Alex |
|
|
2023-10-07 | vanderws |
Process BCSE deep cores - Sidney |
|
|
2023-10-03 | kahmark |
Drone flight for albedo on BCSE and LTAR. Sunny, light breeze. White tarp measure after each. While flying the LTAR, flight lasted unusually long, when it was returning for next battery at 18%, The drone lost complete power about one foot above the home position. Hit the ground and disturbed the mount for the front camera and the ribbon cord that runs from gimbel to the drone system. Repaired. |
|
|
2023-10-03 | vanderws |
G6 ANPP sampling and species composition transects. Three NS transects per plot, main only. Each transect ~2m east of the station flags and numbered 1-3 based on the station number. Stop about every 1m to identify and measure first intercepted plant. ANPP: Station 1 only: 3mS, 0.5mW to NE corner of 1.52x0.66m frame to avoid microplot. All other stations: 3mS, 1.5mW to NE corner of 1.52x0.66m frame. MISGI bagged separately, weeds (if any) sorted to species. - Alyssa, Gloria, Helen, Emily |
|
|
2023-10-01 | vanderws |
Process BCSE deep cores - Sidney |
|
|
2023-09-30 | vanderws |
Process BCSE deep cores - Sidney |
|
|
2023-09-24 | vanderws |
Process BCSE deep cores - Sidney |
|
|
2023-09-23 | vanderws |
Process BCSE deep cores - Sidney |
|
|
2023-09-18 | kahmark |
Removed flags and chambers from Martha’s study on the north baitcrop area BCSE. Additionally, removed all shelters on the Lux Arbor marginal land site. Three reps show high stress. Also, deer pellets in three of the four sheltered areas. ALOT! Only deployed one set of RH/T sensors in rep four this year. |
|
|
2023-09-17 | vanderws |
Process BCSE deep cores - Sidney |
|
|
2023-09-16 | vanderws |
Start processing BCSE 2022 deep soil cores - Sidney |
|
|
2023-09-15 | kahmark |
Lake City Bulk Density cores at all plots 0-4” and 5-9” depths. Shelters moved into winter positions and screws deployed. Checked all loggers G5r2 121 SN 43047, G5r1 127 SN43040, weather SN37075, G5r3 125 SN43043. Checked rain gauge. Some labels were missing, written on core or cap. Possibly sampled a few on an earlier trip with Ashim. Removed RH/T sensors around noon. |
|
|
2023-09-14 | kahmark |
Sampled bulk density at all Escanaba plots 0-4” and 5-9” depths, 2” wide slide hammer sampler AMS, Inc. Checked all loggers G5r1 116 SN42921, G10r3 119 SN43039, G5r2 120 SN43044, 38603 weather station. Removed all shelters by noon. Removed RH/T sensors. Checked rain gage |
|
|
2023-09-13 | kahmark |
Sampled Rhinelander all plots for Bulk Density. 0-4” and 5-9” depth cores (2” wide). Also sampled bulk soil at 22 plots, both depths. Soil was rocky so we used a drain spade shove for this portion. Filled in all holes. Checked data loggers and recorded SNs. SN 38838 weather station, G5r1 126 SN43048, G5r3 124 SN43041, G5r2 118 SN43042. Checked rain gage. Filled in all holes. |
|
|
2023-09-12 | kahmark |
Sampled all plots for bulk density. 0-4” and 5-9”. One inch spacing is due to the non-sampled portion of first core. That one inch is discarded. Checked all loggers and recorded SNs. G5r4 128 SN43046, G5r3 129 SN43045, G5r2 122 SN43049, SN 38839 weather station. |
|
|
2023-09-05 | vanderws |
Started BCSE G4,5,7,9,10 ANPP clipping (finished through G4R2). Station 1: 4mS, 1mW to NE corner of 2 x 0.5m frame, to avoid subplot. Station 2,3: 4mS, 2mW to NE corner of 2 x 0.5m frame. Wasn’t sorted, bagged as UNSRT. - |
|
|
2023-08-30 | dave |
Set fallen HOBO sensor on t-post at Lux Arbor MLE switchgrass site (block 3?). Sensor remained about 1m off ground while collecting previous data (unsure when sensor fell over sideways). Now is upright, approximately 20-30 cm above switchgrass. Inside shelter sensor remains solid. |
|
|
2023-08-08 | kahmark |
Reestablished 8 plots in Martha/Ilyas study with excess N. Kevin hoed and raked each plot to loosen soil. Added Urea 46% after Martha weighed out. Control (0N) and three very high rates of N (400,800,1200gN/m2) in an attempt to reach the threshold of n2O emissions.. Included an image of the addition prior to raking in. On 8/2 Martha weighed out 184,365, and 552g of urea. I checked her calculation on 8/3 and found she did not add the proper amount to the plots. She will adjust today. ADD: 8/7 kept new additions as is. Started sampling 8/8. |
|
|
2023-08-01 | vanderws |
2023 BCSE G8 ANPP/LFLIT
0.5 x .95m wooden frame
North trap… 0mS, 3mE of station 3 to NW corner of frame oriented with short ends on the north/south
South trap… 2mS, 2mE of station 1 to NW corner of frame oriented with short ends on the east/west
Collect newly fallen poplar leaves as LFLIT (if any). 2023 BCSE G8 species composition transects Three NS transects per plot, main only. Each transect ~2m east of the station flags and numbered 1-3 based on the station number. Shift slightly to go down center of two tree rows. Record which tree rows you are between (row 1 is on the west side of the plot). Stop about every 1m to identify and measure first intercepted plant in the understory, not poplar trees.
|
|
|
2023-07-26 | kahmark |
7/26 there was a problem with G5-2 lid solenoid on the trailer system. Found the signal wire to the switch was not connected, but also no lids were firing in close mode. Found a ground wire off as well. Reconnected the G5-2 lid wire and should all be working properly. DW checked the compressor as well. |
|
|
2023-07-20 | kahmark |
Changed battery at weather station Rhinelander. Changed PAR and Solar Radiation sensor at Lake City and Escanaba. Dave installed RH/T and two shelters each at Lake City and Escanaba., Calibration images saved here. |
|
|
2023-07-19 | kahmark |
Rhinelander and Hancock bulk soil sampling KK, DW, AD. Three 0-4” and 4-10” cores per plot. All holes filled. All plots sampled. |
|
|
2023-07-19 | kahmark |
Sampled all plots for bulk soil at Rhinelander and Hancock MLEs. Three cores each of 0-10cm then 10-25cm below. Each core was taken in a one meter square area, one step north of the plot centerline. No more than two meters in the plot (from side) Cores were collected about .5-1 m apart. Holes filled with soil just outside or at the edge of the plot. Took approximately five hours. KK, DW, AD. 2” core samplers. No bulk density. G5 cores taken 7 steps just outside the reserve area. All collected in +N. New battery at the weather station in Rhinelander. We broke two samplers at Rhinelander and stopped bulk soil just prior to finishing R2. All poplars sampled. Used a drain spade to sample at R3 and 4 in Hancock. Created an approximately 4” diameter circle, then inserted spade 10cm and popped up soil. cleared hole, repeated in the 10-25cm depth. Worked very well and was quite fast. Filled in holes when completed. |
|
|
2023-07-18 | kahmark |
Sampled all plots for bulk soil at Escanaba MLE. Three cores each of 0-10cm then 10-25cm below. Each core was taken in a one meter square area, one step north of the plot centerline. No more than two meters in the plot (from side) Cores were collected about .5-1 m apart. Holes filled with soil just outside or at the edge of the plot. Took approximately five hours. KK, DW, AD. 2” core samplers. No bulk density. G5 cores taken 7 steps just outside the reserve area. All collected in +N. Also added two sets of RH/T sensors (one under shelter and one in ambient 0N on a post). Swapped PAR and solar radiation sensors for recalibration. |
|
|
2023-07-18 | kahmark |
Escanaba bulk soil sampling KK, DW, AD. Three 0-4” and 4-10” cores per plot. All holes filled. All plots sampled. |
|
|
2023-07-17 | kahmark |
Sampled all plots for bulk soil at Lake City MLE. Three cores each of 0-10cm then 10-25cm below. Each core was taken in a one meter square area, one step east of the plot centerline. No more than two meters in the plot (from side) Cores were collected about .5-1 m apart. Holes filled with soil just outside or at the edge of the plot. Took approximately five hours. KK, DW, AD. 2” core samplers. No bulk density. G5 cores taken 7 steps just outside the reserve area. All collected in +N. Also added two sets of RH/T sensors (one under shelter and one in ambient 0N on a post). Swapped PAR and solar radiation sensors for recalibration. |
|
|
2023-07-17 | kahmark |
Lake City bulk soil sampling for C and N. Ashim has SURF. Three cores 0-4” and 4-10”, collected 3 per plot of each depth. Bagged and labeled. KK, DW, AD. Very dry conditions. |
|
|
2023-07-10 | kahmark |
Flew BCSE for albedo from 11:30-12:15ish. Fairly windy and difficult to keep the drone steady over the white reference sheet. Sunny |
|
|
2023-07-05 | dave |
Trimmed switchgrass emerging above G5 automated chambers. Found G1-2 lids up, but north clamp down. May have blocked closure of lids over the extended weekend. Fixed north clamp. KK- solenoid for this chamber not currently connected due to issue is either solenoid or clamps. Will address soon. |
|
|
2023-06-27 | dave |
BCSE G1 automated flux chambers removed for fertilization approximately 10am. Returned to service by approximately 12:30pm. |
|
|
2023-06-27 | dave |
BCSE G5, chamber 4 - west lid connector had become unattached. Removed old, rusty connector and replaced with new. (Approximately 2pm.) |
|
|
2023-06-27 | dave |
BCSE G5 automated flux chambers - trimmed all switchgrass in chambers to at/below chamber ceiling. Work done approximately 12:30pm. |
|
|
2023-06-22 | kahmark |
Flew BCSE and LTAR ACSE from 11:00 to 1:00PM. Full sun on the BCSE with reference trp at the end of the flight. At the LTAR site, flew the first set of plots in full sun and the last ten in cloud cover. Then flew the reference tarp in cloud cover. We should use the BCSE full sun reference for the first portion of the LTAR as well. |
|
|
2023-06-01 | kahmark |
Ilya project fertilizer applied today. Control plus three rates of Urea (N), ~8g, ~32g, and ~76g applied over 1m2. Raked in, very dry soil right now (by Martha and Ilya). Plots designed, laid out, and rototilled by Kevin in October 2022 and April 2023 (KK and DW). |
|
|
2023-05-25 | dave |
Moved each G1 automated chamber ~8-12 inches east of previous location to place between corn rows. Planting offset from last year, requiring chamber movement to keep them between rows. All boxes moved between 10am and 11am. |
|
|
2023-05-19 | dave |
3 G1 automated chambers that were moved eastward on 5/12 to accommodate the new sprayer were moved back to previous location (new sprayer didn’t work out, so old sprayer used). Chambers moved approximately 11am. |
|
|
2023-05-05 | kahmark |
Kevin and Dave retilled Ilya’s 2x2m plots in the baitcrop area today. |
|
|
2023-04-14 | kahmark |
Flew Matrice 100 drone for albedo on BCSE around 11AM, took images of T7s for Nameer, all reps. Plan to fly ACSE this afternoon. |
|
|
2023-03-20 | vanderws |
BCSE gas chamber prep - Helen, Emily |
|
|
2023-02-01 | dave |
Shut down compressor in trailer (per Kevin). Changed shut-off valves on manifolds to Swagelok. Checked chambers - all lids in order, all holding pressure except chamber 1-2. |
|
|
2023-01-12 | dave |
eastern lid on chamber 5-5 was found closed on inspection at 8am. Arm attachment had come apart in the usual place. Reattached and arm/lid appear to be working correctly. |
|
|
2023-01-11 | dave |
Chamber lids hyperextending on 1-3,1-5, 5-3, and 5-6. Fixed 1-3, 5-3, and 5-6 - was able to reset and adjust arm/lid attachment to help minimize this in the future. 1-5 reset, but locking nuts could not be loosened to adjust arm. Kevin later adjusted the compressor pressure down to help prevent this from happening going forward. |
|
|
2022-11-21 | kahmark |
Cleared snow from all chambers. removing snow to internal height, clearing area around sampling port, clearing around actuators so they close easily. |
|
|
2022-11-10 | kahmark |
Set up Gelfand carbon vs N2O flux project today. Incorporated 100 lbs of wet weight sorghum into four randomized plots 2m x 2m using a rototiller. Also rototilled the four control plots. Sorghum amended plots have red flag. Four brown paper bags were collected and weighed by Stacey. |
|
|
2022-11-07 | kahmark |
Today, laid out Ilya\’s 2023 project in the north end bait crop to the east of the two track leading to the ball field. (measurements to come). Eight 2m x 2m plots, 2m apart in an area tilled and finished by Joe. The project is 16 m wide east to west, 6m wide north to south and four feet north of the grass alleyway. Previous crop was soybean. Four plots will be rototilled with sorghum at least six inches vertically. Four control plots also rototilled but not amended with carbon. |
|
|
2022-09-29 | dave |
Check on chambers: G1 - 2,3,4,5,6 all had suspected leaf interference. Leaves removed from chamber opening area. G5 - 1,6 had slight leaf/plant interference - cleaned up/removed. Chamber 5 had west arm come loose, so lid half was closed on inspection. Fixed. |
|
|
2022-09-21 | dave |
Flux chamber check: G1 had possible interference from senescing corn leaves on chambers 1, 2, and 6. Leaves cleared to prevent further interference. |
|
|
2022-09-20 | kahmark |
Flew albedo over BCSE at aprox. 1100AM. Hazy sky and no wind. Ht = 25\’ above ground near miscanthus on R2 (highest point in field). Checked trailer instruments and new air conditioning. Running well. Dave checking lids regularly. |
|
|
2022-09-13 | dave |
Checked G1/G5 chambers at BCSE. Removed 3-4 SG plants laying across the face of chamber G5-1. All other chambers okay. |
|
|
2022-09-09 | dave |
Gopher trapped in chamber G5-4. Assisted on rescue/escape. |
|
|
2022-09-07 | kahmark |
Took light saturation curve on switchgrass G5R1 with Mauri. Use for shelter comparison. |
|
|
2022-08-31 | dave |
Flux chamber inspection: G1 - chamber 6 and 3 had 1-2 plants fallen across width of chamber. G5 - chamber 3 had 1-2 plants across width of chamber. Chamber 4 had a whole swath of plants across the chamber - dozens at least. May have been triggered by deer traveling through plot, near chamber. |
|
|
2022-08-30 | kahmark |
Removed all shelters from BCSE between 1-2PM with Dave. All showed under shelter stress. 8/29/2022 removed Lux Arbor shelters (showed some stress). |
|
|
2022-08-19 | vanderws |
BCSE Series 14 gas and soil sampling - Emily, Helen |
|
|
2022-08-19 | dave |
Inspected chambers for malfunctioning lids and plant growth interference. G5 chambers 3, 5, and 6 each had a single strand of switchgrass laying over the open tops, possibly interfering with chamber seal on closure. These were all removed/cleared. |
|
|
2022-08-18 | kahmark |
Lake City MLE drone flight for multispectral and visual (100\’ @ 80% side overlap). Also oblique images for catalog. Under MLE SD Cards file Kevin Docs. Removed shelters, RH/T sensors. Saved camera SD card data and returned (DW). Stress under R1 switchgrass. |
|
|
2022-08-17 | kahmark |
Rhinelander MLE stop. Removed all shelters. Reset cameras, checked asset tags, rain gage, removed RH/T sensors. R1 out did not respond. Drone flight for multispec and visual all. All data from cameras and drone saved on Kahmark computer\documents\MLE SD Cards folder. Shelters out by 11CDT. |
|
|
2022-08-17 | kahmark |
Escanaba stop. Kept shelters in, Todd to remove next week. Removed RH/T sensors from rep 1 and 3 inside outside shelters. Rain gauge okay. Camera cards downloaded and returned. Drone flight for multispec and visual 80\’, 80% side lap (same as others). Camera on G5 was on the ground. Repaired. Repaired gutter on G5R3. |
|
|
2022-08-16 | kahmark |
Hancock visit KK/DW. Took planned multispec and visual drone flight. First south portion done in waypoint mode, shot still every three seconds. Second flight in photomap with shot every two seconds. Checked shelters, kept in place. Some stress under. Will likely take these out later in the season (ask lter techs during anpp transect collection). Future flights give one day notification to Troy, Paul, and Sue. Office now open for signin and signout. |
|
|
2022-07-25 | kahmark |
Added NOx measurements back into the trailer measures. Removed for 10 days to study potential impact of NOx measurements on N2O slope. (3/8/24) Unsure when exactly we created separate loop for NOx vs ICOS as NOx was adding to the N2O in the system. Abide my notes over the 2022 season. We may need to remaove some of the May-July data. |
|
|
2022-07-12 | kahmark |
MLE update. Alleyway mowers cut grass and put hose from shelter gutters under the shelter at some unknown time frame. Will need to look at data and correct for this issue. Todd Martin called this in during his site visit to Hancock.. Manager at Hancock said they mowed three weeks prior to today (or so). |
|
|
2022-06-16 | kahmark |
If you look at the last days trace of N2O from the trailer ICOS, you\’ll see that the lasers never really aligned during startup at 800PM 6/14 which is unusual. You can then see the jump back up at around 1030 today when I realigned manually. Trailer still heating during the day. ICOS having trouble keeping up. . |
|
|
2022-06-14 | kahmark |
If you look at the last days trace of N2O from the trailer ICOS, you\’ll see that the lasers never really aligned during startup at 800PM 6/14 which is unusual. You can then see the jump back up at around 1030 today when I realigned manually. Trailer still heating during the day. ICOS having trouble keeping up. . |
|
|
2022-06-08 | kahmark |
Removed G5 BCSE shelters on 6-7-22 at 3PM, then replaced by noon today. Dave returned G5R1 automated chamber lids by 11AM, turn chambers on in gashog by 1PM. Removed tower yesterday, Cheyenne to return. |
|
|
2022-06-06 | kahmark |
Shelters in on BCSE 5/24 3PM, Lux Arbor 5/27 11 AM, Hancock 5/23, Rhinelander 5/24, Escanaba 5/24, Lake City 5/25. Hancock, Escanaba, Rhinelander, and Lake City have RH/T sensors. Lux and BCSE no sensors available. |
|
|
2022-05-24 | kahmark |
Installed BCSE shelters between 230-330 today. Still need ground screws. No RH/T sensors. |
|
|
2022-05-12 | kahmark |
G1 and G5 chambers now out for planting (G1) and fertilization (G5). Will return this afternoon. First drone flight of the year this afternoon. |
|
|
2022-05-04 | vanderws |
BCSE Series 4 gas and soil (one core from station 1 pooled with one core from station 2) sampling - Kelson, Gloria |
|